OpenStudy (anonymous):

-4x+y=9 rewrite that so y is a function of x

Parth (parthkohli):

y can be expressed as f(x), but first let's isolate y. \( \color{Black}{\Rightarrow y = 4x + 9 }\) \( \color{Black}{\Rightarrow f(x) = 4x + 9 }\) Easy :)

OpenStudy (anonymous):

y=9+4x

OpenStudy (anonymous):

so you flip the y and the 9?

OpenStudy (anonymous):

exactly if you are solving for it you just isolate the variable

OpenStudy (anonymous):

how about this equation?... 1.5y=18-6x

OpenStudy (anonymous):

solve for y?

OpenStudy (anonymous):

wait one minute

OpenStudy (anonymous):

if your solving for y it would be y = 12 -4x

OpenStudy (anonymous):

then find x by using -2, -1, 0, 1